OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (gorv):

y=mx+c m= slope

OpenStudy (gorv):

y=6x+c

OpenStudy (gorv):

for c plug in the value of x and y

OpenStudy (gorv):

-1=6*5+c -1=30+c

OpenStudy (gorv):

c=-31 y=6x-31